# RestoreGrad: Signal Restoration Using Conditional Denoising Diffusion Models with Jointly Learned Prior

### v1

@@ 3 Proposed Method: Integrating DDPM and VAE for Learnable Diffusion Prior
where we approximate the expectations by randomly sampling $(\mathbf{x}_{0},\mathbf{y})\sim q_{\text{data}}(\mathbf{x}_{0},\mathbf{y})$ and $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\text{post}})$, and the summation by sampling $t\sim\mathcal{U}(\{1,\dots,T\})$ (exploiting the independency due to Markov assumption (Nichol & Dhariwal, 2021)) in each training iteration. We also introduce $\eta>0$ for the LR terms and $\lambda>0$ for PM terms, to exert flexible control of the learned latent space.
Sampling of RestoreGrad: In applications that RestoreGrad is mainly concerned with, the target signal $\mathbf{x}_{0}$ is not available in inference time. As in Figure 4, the conditional DDPM then samples $\boldsymbol{\epsilon}\sim p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})=\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\text{prior}})$ from the Prior Net instead; the Posterior Net is no longer needed.
Leveraging Both Target and Conditioner Signal Information for Prior Learning: In the training stage of RestoreGrad, the latent code $\boldsymbol{\epsilon}$ samples from the posterior $q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})$ which exploits both the ground truth signal $\mathbf{x}_{0}$ and conditioner $\mathbf{y}$. It is thus more advantageous than existing works on adaptive priors (e.g., PriorGrad (Lee et al., 2022)) that solely utilize the conditioner $\mathbf{y}$. To observe the benefits brought by the posterior information, we can make the comparison with a variant of RestoreGrad where the Posterior Net is excluded during training, that is,
@@
which basically removes the Posterior Net $\phi$ and only trains the Prior Net $\psi$ and DDPM $\theta$. Interestingly, our results presented later show that RestoreGrad indeed performs better with Posterior Net than without having it in model training.
@@ 4.1.1 Experimental Setup
  • RestoreGrad: We incorporate Prior Net and Posterior Net on top of CDiffuSE. Both modules adopt the ResNet-20 architect (He et al., 2016) suitably modified to 1-D convolutions for waveform processing, each has only 93K learnable parameters (only 2% of the CDiffuSE model).
Configurations: We adopted the basic configurations same as in Lu et al. (2022). The waveforms were processed at 16kHz sampling rate. The number of forward diffusion steps was $T=50$. The variance schedule was $\beta_{t}\in[10^{-4},0.035]$, linearly spaced. The batch size was 16. The fast sampling scheme in Kong et al. (2021) was used in the reverse processes with $S=6$ steps to reduce inference complexity. The inference variance schedule was $\beta^{\text{infer}}_{t}=[10^{-4},10^{-3},0.01,0.05,0.2,0.35]$. Adam optimizer (Kingma & Ba, 2014) was utilized with a learning rate of $2\times 10^{-4}$. We set $\eta=0.1$ and $\lambda=0.5$ for (11). The models were trained on one NVIDIA Tesla V100 GPU (32 GB CUDA memory) and finished 96 epochs in 1 day.
@@ 4.1.2 Results
Improved Model Convergence: As shown in Figure 2 (test set performance), RestoreGrad shows better convergence behavior over PriorGrad (handcrafted prior) and CDiffuSE (standard Gaussian prior). For example, PriorGrad reaches 2.4 in PESQ at 96 epochs, whereas RestoreGrad reaches it in (roughly) 10 epochs, indicating a 10$\times$ speed-up. The results suggest that jointly learning the prior distribution can be beneficial for conditional DDPMs.
Robustness to Reduced Number of Reverse Steps in Inference: RestoreGrad can potentially reduce the inference complexity. In Figure 5, we show how the trained diffusion models tolerate reduction in the number of inference steps. In each model, we trained the network for 96 epochs and then inferenced with $S=3$ reverse steps to compare with the originally adopted $S=6$ steps in Lu et al. (2022). The noise schedule for $S=3$ was $\beta^{\text{infer}}_{t}=[0.05,0.2,0.35]$, a subset of the $S=6$ schedule that resulted in best performance. We can see that the baseline DDPM is most sensitive to the step reduction, while PriorGrad shows certain tolerance as leveraging a closer-to-data prior distribution. Finally, RestoreGrad barely degrades with reduced sampling steps, echoing that a better prior has been obtained as it recovers higher fidelity signal even in fewer reverse steps.
@@
Comparison to Fully-Trained CDiffuSE: We present in Table 4.1.2 more detailed comparison of RestoreGrad with the baseline CDiffuSE. Here, the scores of CDiffuSE were directly taken from the results reported in Lu et al. (2022) where the model has been fully trained for 445 epochs. For PriorGrad and RestoreGrad we report the mean±$\pm$±std computed based on results of 10 independent samplings. We can see that with RestoreGrad applied, the SE model can achieve better performance over the baseline CDiffuSE by only training for 96 epochs (4.6 times lesser than the baseline) in all the metrics. In addition, halving the number of reverse steps in inference still maintains better performance than the fully-trained CDiffuSE and also the PriorGrad.
Signal Quality and Encoder Size Trade-Offs / Complexity Analysis: We further present results using three different model sizes (24K, 93K, 370K) for the Prior and Posterior Nets (i.e., encoders) in Table 4.1.2, along with latency and GPU memory usage. The results clearly show that the restored speech quality improves with an increasing encoder size. This indicates that there is a trade-off between the restoration signal quality and encoder model complexity. Notably, the latency and memory usage of the encoder modules are relatively small compared to the DDPM processing (decoding), suggesting that RestoreGrad is capable of achieving improved performance without incurring considerable increase in complexity compared to the adopted DDPM model.
@@
Table 2: SE comparison of RestoreGrad models using encoder modules of different sizes and the corresponding latency and GPU memory usage (measured on one NVIDIA Tesla V100 GPU).
{NiceTabular}
@@
(# params) DDPM Encoder DDPM Encoder
Tiny (24K) 2.48 3.11 5.10 13.74 229.5 4.5 74.4M 4.8M
Base (93K) 2.51 3.14 5.92 14.74 230.1 5.1 74.5M 7.7M
Large (370K) 2.54 3.16 6.15 15.01 231.7 6.0 74.7M 13.6M
@@ 4.2.1 Experimental Setup
Dataset: Following Özdenizci & Legenstein (2023), we consider the IR task of recovering clean images from their degraded versions contaminated by synthesized noises corresponding to different weather conditions. Two datasets are considered, where one is a weather-specific dataset called RainDrop (Qian et al., 2018) and the other is a multi-weather dataset named AllWeather (Valanarasu et al., 2022). The RainDrop dataset consists of images captured with raindrops on the camera sensor which obstruct the view. It has 861 training images with synthetic raindrops, and a test set of 58 images dedicated for quantitative evaluations. The AllWeather dataset is a curated training dataset from Valanarasu et al. (2022), which has 18,069 samples composed of subsets of training images from Snow100K (Liu et al., 2018), Outdoor-Rain (Li et al., 2019) and RainDrop (Qian et al., 2018), in order to create a balanced training set across three weather conditions.
@@
  • RestoreGrad: We incorporate the encoder modules, Prior Net and Posterior Net, on top of the baseline DDPM. Both encoder modules adopt the ResNet-20 architect (He et al., 2016) with only 0.27M learnable parameters, significantly smaller ($<0.3\%$) than the baseline DDPM model.
Configurations: We used Adam optimizer with a learning rate of $2\times 10^{-5}$. An exponential moving average with a weight of 0.999 was applied. We used $T=1000$ and linear noise schedule $\beta_{t}\in[10^{-4},0.02]$, same as Özdenizci & Legenstein (2023). A batch size of 4 was used. The models were trained on two NVIDIA Tesla V100 GPUs of 32 GB CUDA memory and finished training for 9,261 epochs on the RainDrop dataset in 12 days and 887 epochs on the AllWeather dataset in 21 days.
4.2.2 Results
Model Convergence: As presented in Figure 2 (test set performance), RestoreGrad demonstrates faster convergence and better restored image quality over the baseline DDPM (RainDropDiff). For example, RainDropDiff reaches 0.143 in LPIPS at 9.2k epochs, while RestoreGrad reaches it in 1.8k epochs only, indicating a 5$\times$ speed-up due to the effectiveness of the prior learning scheme.
Comparison with Existing IR Models: We compare our method on multi-weather cases in Table 4.2.2 with All-in-One (Li et al., 2020) and TransWeather (Valanarasu et al., 2022), where the models were trained on the AllWeather dataset and tested on the three weather-specific test sets. The numbers of the compared models were taken from Özdenizci & Legenstein (2023), where the WeatherDiff was trained for 1,775 epochs and inferenced with $S=25$ steps. Our RestoreGrad was trained for only 887 epochs (2$\times$ fewer than WeatherDiff) and inferenced with $S=10$ steps to already achieve the best performance in almost all test schemes.
@@ 4.2.2 Results
WS RESCAN (Li et al., 2018) 26.08 0.8108 HRGAN (Li et al., 2019) 21.56 0.8550 RaindropAttn (Quan et al., 2019) 31.44 0.9263
DesnowNet (Liu et al., 2018) 27.17 0.8983 PCNet (Jiang et al., 2021) 26.19 0.9015 AttentiveGAN (Qian et al., 2018) 31.59 0.9170
@@
Other IR Tasks: Our method demonstrates the advantages of adopting learnable priors also in image deblurring and super-resolution tasks, suggesting its generality. We refer the reader to Appendix D.3 for the results and discussion.
@@ 5 Related Work
Diffusion-Based Signal Restoration: Built on top of the diffusion models for audio generation, e.g., Kong et al. (2021); Chen et al. (2020); Leng et al. (2022), many SE models have been proposed. The pioneering work of Lu et al. (2022) introduced conditional DDPMs to the SE task and demonstrated the potential. Other works (Serrà et al., 2022; Welker et al., 2022; Richter et al., 2023; Yen et al., 2023; Lemercier et al., 2023; Tai et al., 2023a) have also attempted to improve SE by exploiting diffusion models. In the vision domain, diffusion models have demonstrated impressive performance for IR tasks (Li et al., 2023; Zhu et al., 2023; Huang et al., 2024; Luo et al., 2023; Xia et al., 2023; Fei et al., 2023; Hurault et al., 2022; Liu et al., 2023a; Chung et al., 2023b, a; Zhou et al., 2024; Xiao et al., 2024; Zheng et al., 2024). A notable IR work is (Özdenizci & Legenstein, 2023) that achieved impressive performance on several benchmark datasets for restoring vision in adverse weather conditions. Despite showing promising results, existing works have not fully exploited prior information about the data as they mostly settle on standard Gaussian priors.
@@ 6 Conclusion
We investigated the potential of jointly learning the prior distribution with the conditional DDPM for improved efficiency. We demonstrated the advantages of leveraging learning-based priors with RestoreGrad, providing a more systematic way of estimating the prior than existing selections. A limitation of the current work is that it focuses on signal restoration applications, where we suitably assume a zero-mean Gaussian prior and only learn its covariance. In the future, it can be interesting to explore using a more generic prior form and extend the idea to other applications.
@@
Proposition 3.1 (Incorporation of diffusion process into VAE). By introducing a sequence of hidden variables $\mathbf{x}_{1:T}$, under the setup of conditional diffusion models where the Markov Chain assumption is employed on the forward process $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})\coloneq\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ and the reverse process $p_{\theta}(\mathbf{x}_{0:T}|\mathbf{y})\coloneq p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ parameterized by a DDPM $\theta$, and assuming that $\mathbf{x}_{T}=\boldsymbol{\epsilon}$ (i.e., the latent noise of DDPM samples from the VAE latent distribution), we have the lower bound on $\log p(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})$ in the reconstruction term of the VAE as:
$\log p_{\theta}(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})\geq\mathbb{E}_{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\left[\log\frac{p_{\theta}({\mathbf{x}_{0:T}}|\mathbf{y})}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right],$
which is the ELBO of the conditional DDPM.
@@
We lower bound the conditional data log-likelihood $\log p_{\theta}(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})$ by utilizing a sequence of hidden latent representations $\mathbf{x}_{1:T}$ and the approximate variational distribution $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$:
@@
Proposition 3.2 (RestoreGrad). Assume the prior and posterior distributions are both zero-mean Gaussian, parameterized as $p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\boldsymbol{\Sigma}_{\text{prior}}(\mathbf{y};\psi))$ and $q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\boldsymbol{\Sigma}_{\text{post}}(\mathbf{x}_{0},\mathbf{y};\phi))$, respectively, where the covariances are estimated by the Prior Net $\psi$ (taking $\mathbf{y}$ as input) and Posterior Net $\phi$ (taking both $\mathbf{x}_{0}$ and $\mathbf{y}$ as input). Let us simply use $\boldsymbol{\Sigma}_{\text{prior}}$ and $\boldsymbol{\Sigma}_{\text{post}}$ hereafter to refer to $\boldsymbol{\Sigma}_{\text{prior}}(\mathbf{y};\psi)$ and $\boldsymbol{\Sigma}_{\text{post}}(\mathbf{x}_{0},\mathbf{y};\phi)$ for concise notation. Then, with the direct sampling property in the forward path $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ at arbitrary timestep $t$ where $\boldsymbol{\epsilon}\sim q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})$, and assuming the reverse process has the same covariance as the true forward process posterior conditioned on $\mathbf{x}_{0}$, by utilizing the conditional DDPM $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$ as the noise estimator of the true noise $\boldsymbol{\epsilon}$, we have the modified ELBO associated with (9):
$-ELBO=$ $\underbrace{\frac{\bar{\alpha}_{T}}{2}\mathbb{E}_{\mathbf{x}_{0}}\lVert\mathbf{x}_{0}\rVert^{2}_{\boldsymbol{\Sigma}^{-1}_{\text{post}}}+\frac{1}{2}\log\lvert\boldsymbol{\Sigma}_{\text{post}}\rvert}_{\text{Latent Regularization (LR) terms}}+\underbrace{\sum_{t=1}^{T}\gamma_{t}\mathbb{E}_{(\mathbf{x}_{0},\mathbf{y}),\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\text{post}})}\lVert\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)\rVert^{2}_{\boldsymbol{\Sigma}^{-1}_{\text{post}}}}_{\text{Denoising Matching (DM) terms}}$
@@
The --ELBO can be shown to be expanded as (Luo, 2022; Sohl-Dickstein et al., 2015; Ho et al., 2020):

### v3

@@ 3.1 Training of RestoreGrad
where we approximate the expectations by randomly sampling $(\mathbf{x}_{0},\mathbf{y})\sim q_{\text{data}}(\mathbf{x}_{0},\mathbf{y})$ and $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\text{post}})$, and the summation over $t$ by sampling $t\sim\mathcal{U}(\{1,\dots,T\})$ (exploiting the independency due to Markov assumption (Nichol & Dhariwal, 2021)) in each training iteration. We have also introduced $\eta>0$ for the LR terms and $\lambda>0$ for PM terms, to exert flexible control of the learned latent space.
3.2 Sampling of RestoreGrad
In applications that RestoreGrad is mainly concerned with, the target signal $\mathbf{x}_{0}$ is not available in inference time. As in Figure 4, the conditional DDPM then samples $\boldsymbol{\epsilon}\sim p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})=\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\text{prior}})$ from the Prior Net instead; the Posterior Net is no longer needed.
3.3 The Role of Posterior Information
In the training stage of RestoreGrad, the latent code $\boldsymbol{\epsilon}$ samples from the posterior $q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})$ which exploits both the ground truth signal $\mathbf{x}_{0}$ and conditioner $\mathbf{y}$. It is thus more advantageous than existing works on adaptive priors (e.g., PriorGrad (Lee et al., 2022)) that only utilize the conditioner $\mathbf{y}$. To observe the benefits brought by the posterior information, we can make comparison with a variant of RestoreGrad where the Posterior Net is excluded in training:
@@ 3.3 The Role of Posterior Information
which basically removes the Posterior Net $\phi$ and only trains the Prior Net $\psi$ and DDPM $\theta$. Interestingly, our experimental results show that RestoreGrad indeed performs better with the Posterior Net than without it in model training.
@@ 4.1.1 Experimental Setup
  • RestoreGrad: We incorporate Prior Net and Posterior Net on top of CDiffuSE. Both modules adopt the ResNet-20 architect (He et al., 2016), suitably modified to 1-D convolutions for waveform processing, each with only 93K learnable parameters (only 2% of the CDiffuSE model).
Configurations: We adopted the basic configurations same as in Lu et al. (2022). The waveforms were processed at 16kHz sampling rate. The number of forward diffusion steps was $T=50$. The variance schedule was $\beta_{t}\in[10^{-4},0.035]$, linearly spaced. The batch size was 16. The fast sampling scheme in Kong et al. (2021) was used in the reverse process with 6 steps to reduce inference complexity, with the 6-step inference variance schedule $\beta^{\text{infer}}_{t}=[10^{-4},10^{-3},0.01,0.05,0.2,0.35]$. Adam optimizer (Kingma & Ba, 2014) was utilized with a learning rate of $2\times 10^{-4}$. We set $\eta=0.1$ and $\lambda=0.5$ for (11).
@@ 4.1.2 Results
Improved Model Convergence: As shown in Figure 2 (test set performance), RestoreGrad shows better convergence behavior over PriorGrad (using handcrafted prior) and CDiffuSE (using standard Gaussian prior). For example, PriorGrad reaches 2.4 in PESQ at 96 epochs, whereas RestoreGrad reaches it in (roughly) 10 epochs, indicating a 10$\times$ speed-up. The results suggest that jointly learning the prior distribution can be beneficial for conditional DDPMs.
Robustness to Reduced Reverse Steps in Inference: RestoreGrad can potentially reduce the inference complexity. In Figure 5, we show how the trained diffusion models withstand the reduction in the number of inference steps. In each model, we trained the network for 96 epochs and then inferenced with 3 reverse steps to compare with the originally adopted 6-step scheme in Lu et al. (2022). The noise schedule for the 3-step scheme was $\beta^{\text{infer}}_{t}=[0.05,0.2,0.35]$, a subset of the 6-step schedule that resulted in best performance. We can see that the baseline DDPM is most sensitive to the step reduction, while PriorGrad shows certain resistance as leveraging a closer-to-data prior distribution. Finally, RestoreGrad barely degrades with reduced sampling steps, echoing that a better prior has been obtained as it recovers higher fidelity signal even in fewer reverse steps.
@@
Comparison to Fully-Trained CDiffuSE: We present in Table 4.1.2 more detailed comparison of RestoreGrad with the baseline CDiffuSE. Here, the scores of CDiffuSE were directly taken from the results reported in Lu et al. (2022) where the model has been fully trained for 445 epochs. For PriorGrad and RestoreGrad we report the mean±$\pm$±std computed based on results of 10 independent samplings. We can see that with RestoreGrad applied, the SE model can achieve better performance over the baseline CDiffuSE by only training for 96 epochs (4.6 times lesser) in all the metrics. In addition, halving the number of reverse steps in inference time still maintains better performance than the fully-trained CDiffuSE and also the PriorGrad.
@@
Signal Quality and Encoder Complexity Trade-Offs: We further present results using three different model sizes (24K, 93K, 370K) for the Prior and Posterior Nets (encoders) in Table 2, along with latency and GPU memory usage (presented as the ratio of encoder to DDPM). The results clearly show that the restored speech quality improves with increasing encoder size. This indicates that there is a trade-off between the restoration signal quality and encoder model complexity. Notably, the latency and memory usage of the encoder modules are relatively small compared to the DDPM decoding ($<$2.6% latency and $<$18.2% memory usage of the DDPM processing), suggesting that RestoreGrad is capable of achieving improved performance without incurring considerable increase in complexity compared to the adopted DDPM model.
Table 2: SE comparison of RestoreGrad models using encoder modules of different sizes and the corresponding latency and GPU memory usage (measured on one NVIDIA Tesla V100 GPU) presented as the ratio of encoder to DDPM.
@@
Tiny (24K) 2.48 3.11 5.10 13.74 1.9% 6.5%
Base (93K) 2.51 3.14 5.92 14.74 2.2% 10.3%
Large (370K) 2.54 3.16 6.15 15.01 2.6% 18.2%
@@ 4.2.1 Experimental Setup
Dataset: Following Özdenizci & Legenstein (2023), we consider the IR task of recovering clean images from their degraded versions contaminated by synthesized noises corresponding to different weather conditions. Two datasets are considered, where one is a weather-specific dataset called RainDrop (Qian et al., 2018) and the other is a multi-weather dataset named AllWeather (Valanarasu et al., 2022). The RainDrop dataset consists of images captured with raindrops on the camera sensor which obstruct the view. It has 861 training images and a test set of 58 images dedicated for quantitative evaluations. The AllWeather dataset is a curated training dataset from Valanarasu et al. (2022), which has 18,069 samples composed of subsets of training images from Snow100K (Liu et al., 2018), Outdoor-Rain (Li et al., 2019) and RainDrop (Qian et al., 2018), in order to create a balanced training set across three weather conditions.
@@
  • RestoreGrad: We incorporate the encoder modules, Prior Net and Posterior Net, on top of the baseline DDPMs. Both encoder modules adopt the ResNet-20 architect (He et al., 2016) with only 0.27M learnable parameters, significantly smaller ($<0.3\%$) than the baseline DDPM models.
Configurations: We used Adam optimizer with a learning rate of $2\times 10^{-5}$. An exponential moving average with a weight of 0.999 was applied. We used $T=1000$ and linear noise schedule $\beta_{t}\in[10^{-4},0.02]$, same as Özdenizci & Legenstein (2023). A batch size of 4 was used.
@@
WS RESCAN (Li et al., 2018) 26.08 0.8108 HRGAN (Li et al., 2019) 21.56 0.8550 AttentiveGAN (Qian et al., 2018) 31.59 0.9170
DesnowNet (Liu et al., 2018) 27.17 0.8983 PCNet (Jiang et al., 2021) 26.19 0.9015 RaindropAttn (Quan et al., 2019) 31.44 0.9263
@@
+ RestoreGrad (ours) – trained longer 31.16 0.9175 + RestoreGrad (ours) – trained longer 30.70 0.9418 + RestoreGrad (ours) – trained longer 32.26 0.9414
@@
4.2.2 Results
Model Convergence: As presented in Figure 2 (test set performance), RestoreGrad demonstrates faster convergence and better restored image quality over the baseline DDPM (RainDropDiff). For example, RainDropDiff reaches 0.143 in LPIPS at 9.2K epochs, while RestoreGrad reaches it in 1.8K epochs only, indicating a 5$\times$ speed-up due to the effectiveness of the joint prior learning scheme.
Comparison with Existing IR Models: We compare our method with existing models of the multi-weather (MW) type in Table 4.2.1, including All-in-One (Li et al., 2020) and TransWeather (Valanarasu et al., 2022) in addition to WeatherDiff (Özdenizci & Legenstein, 2023). The models were all trained on the AllWeather dataset and the performance numbers were taken from Özdenizci & Legenstein (2023), where the WeatherDiff was trained for 1,775 epochs and inferenced with 25 steps. Our RestoreGrad was trained for only 887 epochs (2$\times$ fewer) and inferenced with 10 steps to already achieve the best performance in the MW category. In addition, when trained for more epochs (1,551 epochs, which is 1.14$\times$ fewer than WeatherDiff), RestoreGrad achieves further improvements as shown in the last row the table. On the other hand, our method, while trained on multi-weather data, achieves a comparable performance to the recently proposed Diffusion Texture Prior Model (DTPM) (Ye et al., 2024) individually trained on the three weather-specific (WS) datasets. Notably, our method does not require a pretraining stage on a large dataset of high-quality images, unlike DTPM which is pretrained on 55,000 images samples. This suggests the generality and effectiveness of our method to improve baseline DDPM models.
Other IR Tasks: Our method demonstrates the advantages of learnable priors in image deblurring and super-resolution tasks, suggesting its generality. See Appendix D.2.
4.3 Generalization to Out-of-Distribution (OOD) and Realistic Data
We have so far evaluated the models on in-domain scenarios with synthetic noisy data where RestoreGrad has shown substantial improvements. A natural question is whether the demonstrated improvements have actually come at the expense of the model’s generalizability to unseen or realistic data. To address the concern, we evaluate the IR models on two additional datasets,RainDS-Real from Quan et al. (2021) and Snow100K-Real from Liu et al. (2018) consisting of real-world images, using the reference-free Natural Image Quality Evaluator (NIQE) metric (Mittal et al., 2012) (a lower score indicates better quality). In Table 4.3 we see that RestoreGrad is able to perform on par with or better than WeatherDiff and the non-generative model of TransWeather. For OOD testing, we evaluate the SE models on the CHiME-3 dataset (Barker et al., 2017) unseen during model training. Table 4.3 compares RestoreGrad with CDiffuSE also trained for 96 epochs, DOSE (Tai et al., 2023a), and two discriminative SE models, Demucs (Defossez et al., 2020) and WaveCRN (Hsieh et al., 2020). We can see that RestoreGrad is able to perform equally well as the CDiffuSE while outperforming DOSE and the non-generative SE models. The results show that RestoreGrad is capable of improving in-domain performance while maintaining desirable generalizability of generative models.
Table 5: Evaluation on realistic image datasets of the IR models trained on synthetic images of AllWeather training set.
{NiceTabular}
lccc Methods Gen. RainDS-Real Snow-Real
NIQE $\downarrow$ NIQE $\downarrow$
TransWeather (Valanarasu et al., 2022) N 4.005 3.161
WeatherDiff (Özdenizci & Legenstein, 2023) Y 3.050 2.985
+ RestoreGrad (ours) Y 2.556 3.015
{NiceTabular}
[…]
Unprocessed - 1.27 2.61 1.93 1.88 7.51
Demucs (Defossez et al., 2020) N 1.38 2.50 2.08 1.88 -
WaveCRN (Hsieh et al., 2020) N 1.43 2.53 2.03 1.91 -
DOSE (Tai et al., 2023a) Y 1.52 2.71 2.15 2.06 -
CDiffuSE (Lu et al., 2022) Y 1.55 2.87 2.09 2.15 7.67
+ RestoreGrad (ours) Y 1.54 2.88 2.14 2.16 8.45
@@ 5 Related Work
Diffusion-Based Signal Restoration: Built on top of the diffusion models for audio generation, e.g., Kong et al. (2021); Chen et al. (2020); Leng et al. (2022), many SE models have been proposed. The pioneering work of Lu et al. (2022) introduced conditional DDPMs to the SE task and demonstrated the potential. Other works (Serrà et al., 2022; Welker et al., 2022; Richter et al., 2023; Yen et al., 2023; Lemercier et al., 2023; Tai et al., 2023a) have also attempted to improve SE by exploiting diffusion models. In the vision domain, diffusion models have demonstrated impressive performance for IR tasks (Li et al., 2023; Zhu et al., 2023; Huang et al., 2024; Luo et al., 2023; Xia et al., 2023; Fei et al., 2023; Hurault et al., 2022; Liu et al., 2023a; Chung et al., 2023b, a; Zhou et al., 2024; Xiao et al., 2024; Zheng et al., 2025; Liu et al., 2024; Ye et al., 2024), and a comprehensive view of recent advancements is provided by He et al. (2025). A notable IR work is by Özdenizci & Legenstein (2023) that achieved impressive performance on several benchmark datasets for restoring vision in adverse weather conditions. Despite showing promising results, existing works have not fully exploited prior information about the data as they mostly settle on standard Gaussian priors.
@@ 6 Conclusion
We investigated the potential of learning the prior distribution in tandem with the conditional DDPM for improved efficiency. We demonstrated the advantages of RestoreGrad that leverages learning-based priors, providing a more systematic way of estimating the prior than existing selections. A limitation of the current work is that it focuses on signal restoration applications, where we suitably assume a zero-mean Gaussian prior and only learn its covariance. In the future, it can be interesting to explore using a more generic prior form and extend the idea to other applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
Proposition 3.1 (Incorporation of diffusion process into VAE). By introducing a sequence of hidden variables $\mathbf{x}_{1:T}$, under the setup of conditional diffusion models where the Markov Chain assumption is employed on the forward process $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})\coloneq\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ and the reverse process $p_{\theta}(\mathbf{x}_{0:T}|\mathbf{y})\coloneq p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ parameterized by a DDPM $\theta$, and assuming that $\mathbf{x}_{T}=\boldsymbol{\epsilon}$ (i.e., the latent noise of DDPM samples from the VAE latent distribution), we have the lower bound on $\log p(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})$ in the reconstruction term of the VAE (7) as:
$\log p(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})\geq\mathbb{E}_{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\left[\log\frac{p_{\theta}({\mathbf{x}_{0:T}}|\mathbf{y})}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right],$
which is the ELBO of the conditional DDPM (i.e., the conditional version of (3)).
@@
We lower bound the conditional data log-likelihood $\log p(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})$ by utilizing a sequence of hidden latent representations $\mathbf{x}_{1:T}$ and the approximate variational distribution $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$:
@@
Proposition 3.2 (RestoreGrad). Assume the prior and posterior distributions are both zero-mean Gaussian, parameterized as $p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\boldsymbol{\Sigma}_{\text{prior}}(\mathbf{y};\psi))$ and $q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\boldsymbol{\Sigma}_{\text{post}}(\mathbf{x}_{0},\mathbf{y};\phi))$, respectively, where the covariances are estimated by the Prior Net $\psi$ (taking $\mathbf{y}$ as input) and Posterior Net $\phi$ (taking both $\mathbf{x}_{0}$ and $\mathbf{y}$ as input). Let us simply use $\boldsymbol{\Sigma}_{\text{prior}}$ and $\boldsymbol{\Sigma}_{\text{post}}$ hereafter to refer to $\boldsymbol{\Sigma}_{\text{prior}}(\mathbf{y};\psi)$ and $\boldsymbol{\Sigma}_{\text{post}}(\mathbf{x}_{0},\mathbf{y};\phi)$ for concise notation. Then, with the direct sampling property in the forward path $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ at arbitrary timestep $t$ where $\boldsymbol{\epsilon}\sim q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})$, and assuming the reverse process has the same covariance as the true forward process posterior conditioned on $\mathbf{x}_{0}$, by utilizing the conditional DDPM $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$ as the noise estimator of the true noise $\boldsymbol{\epsilon}$, we have the modified ELBO, $-\mathcal{L}(\theta,\phi,\psi)$, associated with (9):
$\mathcal{L}(\theta,\phi,\psi)=$ $\underbrace{\frac{\bar{\alpha}_{T}}{2}\mathbb{E}_{\mathbf{x}_{0}}\lVert\mathbf{x}_{0}\rVert^{2}_{\boldsymbol{\Sigma}^{-1}_{\text{post}}}+\frac{1}{2}\log\lvert\boldsymbol{\Sigma}_{\text{post}}\rvert}_{\text{Latent Regularization (LR) terms}}+\underbrace{\sum_{t=1}^{T}\gamma_{t}\mathbb{E}_{(\mathbf{x}_{0},\mathbf{y}),\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\text{post}})}\lVert\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)\rVert^{2}_{\boldsymbol{\Sigma}^{-1}_{\text{post}}}}_{\text{Denoising Matching (DM) terms}}$
@@
The --ELBO $\mathcal{L}(\theta,\phi)$ can be shown to be expanded as (Luo, 2022; Sohl-Dickstein et al., 2015; Ho et al., 2020):